OpenStudy (anonymous):

A rectangle has sides 60 cm long and 220 cm wide, and a square has sides 80 cm long. Find the ratio of (a) perimeters and (b) the areas

OpenStudy (mathstudent55):

Perimeter and Area of Rectangle: PR = 2L + 2W AR = LW (L = length, W = width) Perimeter and Area of Square: PS = 4S AS = S^2 (S = length of side of square) You are looking for: PR/PS and AR/AS

OpenStudy (anonymous):

A

OpenStudy (anonymous):

= 7:4 and B = 33:16 am i right?

OpenStudy (mathstudent55):

Yes, correct!